# Scaling the Variational Quantum Eigensolver for Dynamic Portfolio Optimization

### v1

@@ Abstract
This work explores the potential of the Variational Quantum Eigensolver in solving Dynamic Portfolio Optimization problems surpassing the 100 qubit utility frontier. We systematically analyze how to scale this strategy in complexity and size, from 6 to 112 qubits, by testing different combinations of ansatz and optimizer on a real Quantum Processing Unit. We achieve best results when using a combination of a Differential Evolution classical optimizer and an ansatz circuit tailored to both the problem and the properties of the Quantum Processing Unit.
@@ I Introduction
Figure 1: Optimization cost distributions obtained for the starting VQE approach (orange line), and the fine-tuned VQE implementation (red line) that we develop in this work. These distributions are obtained from QPU calculations in the IBM Torino solving the 112 qubits (XXL sized) DPO problem presented in Section II. For reference we also give the solution obtained using the Gurobi classical optimizer (black dashed line) and the offset of the problem (average of the random distribution, gray dashed line). We add the optimization costs labels “better” and “worse” to reinforce that we address the DPO as a minimization problem. To obtain each cost distribution we account for how many times we found investment strategies (quantum states) with the same associated cost (rounded to the second decimal value).
@@ II.1 Formulation of the Dynamic Portfolio Optimization problem
$F(\bm{\Omega}),R(\bm{\Omega})$ and $C(\bm{\Omega})$, in Eq. (2) are calculated over a range of $N_{a}$ assets and across $N_{t}$ time steps. In the expression above, $\gamma$ is the risk aversion coefficient. We found that (for the DPO problems solved here) $\gamma=1000$ ensures same order of magnitude for the return and the risk in Eq. (2) [^1] In our work we chose parameter values that ensure all key aspects of the objective function in Eq. (2) are well represented, in order to capture the full complexity of the problem. For example, we do not chose the value of the risk aversion coefficient $\gamma$ according to industry standards, but we chose a $\gamma$ value that ensures the risk function $R(\bm{\Omega})$ and the expected return function $F(\bm{\Omega})$ have the same influence on the final solution.. Specifically, the expected return is defined as
@@
where $\bm{\mu}_{t}=(\mu_{t,0},\dots,\mu_{t,N_{a}-1})$ is the return vector. The components of $\bm{\mu}_{t}$ are given by the logarithmic return  [^2] The expression of the logarithmic expected return we use in Eq. (4) can be traced back to the expected bare return of the investment. Specifically, the $\omega_{a,t}$ money invested in an asset $a$ at time $t$ becomes $\mu^{\text{bare}}_{t,a}\times\omega_{t}$ at the next investment window, with the bare return $\mu^{\text{bare}}_{t,a}=1+(P_{t+1,a}-P_{t,a})/P_{t,a}$. In Eq. (4) we take the logarithmic of the bare return, $\mu_{t,a}=\log(\mu^{\text{bare}}_{t,a})$. of each asset $a$, which are calculated from the changes of their prices at time $t+1$, $P_{t+1,a}$, compared to the previous time price, $P_{t,a}$,
@@
Note that for $t=0$ we have a term $\bm{\omega}_{-1}$ in Eq. (6), which corresponds to the initial condition of the problem. We chose to set $\bm{\omega}_{-1}=\bm{0}_{N_{a}}$ as if there is no initial investment.
@@ II.2 Dynamic Portfolio Optimization problem as a Quadratic Unconstrained Binary Optimization problem
In order to work with the VQE algorithm, we must transform the objective function $O(\bm{\Omega})$ in Eq. (2) into a QUBO problem. By definition, a QUBO problem consists of binary variables with terms that are either linear or quadratic, and it must exclude constraints.
@@
$\tilde{\omega}_{a,t}:=\frac{\omega_{a,t}}{K}.$ (10)
@@
Regarding the QUBO formulation, we first tackle the transaction cost function. Due to the absolute value introduced within the transaction costs in Eq. (6), the function is neither quadratic nor linear, so we perform a quadratic approximation of the absolute value, which leads us to the expression
@@
with $x_{t,a,r}$ binary variables, and the subindices $t$ and $a$ representing the time and the asset of the investment, respectively. The subindex $r$ ranges from 00 to $N_{r}-1$, where $N_{r}$ is the amount of binary variables used to describe the normalized investment ($\tilde{\omega}_{a,t}$) for each asset at each time (increasing $N_{r}$ enhances the precision level defining such investment). Note that, with this conversion, the maximum money invested per asset ($K^{\prime}$ in Eq. (11)) is $K^{\prime}=2^{N_{r}}-1$.
For convenience, we convert the QUBO problem in Eq. (15) into an Ising Hamiltonian, $H_{\text{Ising}}$, since most quantum optimizers (e.g., quantum annealing) focus on finding the minimum eigenvalue of such Hamiltonians (instead of directly operating on the QUBO problem). Thus, we perform a change of variables from the binary variables $x_{t,a,r}$ to Pauli operators $Z_{q}$,
@@ III.1 Classical Benchmarks
Table 3: Sharpe ratio results of the investment strategies associated to each minimum optimization cost values shown in Table 2.
@@ III.2 Quantum-Inspired Benchmark
In addition to our classical benchmarks, we develop a quantum-inspired Tensor Network algorithm using the Python library QUIMB [39]. We implement a simulated adiabatic evolution (SAE) [40] algorithm, following Ref. [41]. The SAE allows us to have a benchmark for a quantum strategy run on classical hardware. The details on the implementation of the SAE are beyond the scope of the main text, and are described on Appendix F.
@@ IV Variational Quantum Eigensolver
We next implement the VQE algorithm to solve the QUBO problem formulated in Eq. (15) in real QPU devices. The VQE algorithm is a leading approach in solving quantum optimization problems, demonstrating significant success on current NISQ-era devices [42, 43, 44]. In this section, we review our implementation of VQE following the scheme in Fig. 3. Specifically our implementation consist on five steps:
Figure 3: Scheme of the VQE optimization process (described in Section IV): Steps 1-3 comprise generating an ansatz circuit for a given set of parameters $\bm{\theta}$ (step 1), measuring the expected value of the Ising Hamiltonian $H_{\text{Ising}}$ using the estimator primitive of IBM quantum services (step 2), and optimizing $\bm{\theta}$ with a classical computer (step 3). If the optimal $\bm{\theta}$ is not found in step 3, the process returns to step 1. Once the optimal $\bm{\theta}$ is found, the algorithm proceeds to sample the solution (steps 4 and 5), which involves sampling the ansatz circuit for the optimal $\bm{\theta}$ (step 4, where we use the sampler primitive of IBM quantum services) and identifying the optimal sampled state (step 5).
@@
    Finally, we compute the optimization cost associated with each bit string, $c(b)=\braket{b}{\hat{H}}{b}$, and identify the solution bit string |$\ket{b_{\text{s}}}$ that results in the smallest value of the optimization cost.
@@ V Quantum Dynamic Portfolio Optimization
As mentioned in the previous section, the performance of the VQE algorithm relies on the chosen classical optimizer and ansatz. In this section, we explore this dependence as we scale the DPO problem from a size requiring 6 qubits to a large scale problem requiring 112 qubits (see Table 1) and we systematically solve each problem size for a better comprehension of the scaling for each method. Upscaling the problem not only increases the number of qubits but also the depth of the circuits, leading to decoherence and gate-associated quantum noise. We remark that in our work we do not use any quantum error mitigation. Thus, we aim to find VQE implementations that are robust enough to quantum noise such that their raw implementation gives satisfactory results.
@@
It is important to note that every circuit must be transpiled [49] before running it on QPU. This transpilation primarily depends on the basic gates of the quantum computer and the arrangement of its physical qubits. For example, if two logical qubits that need to be connected are not directly linked on the QPU, SWAP gates are introduced to achieve this entanglement. The transpilation process can significantly increase the depth of each circuit. In this work, we discuss depth after transpilation.
The comparison between methods for all problem sizes in the subsections bellow is built upon the results in Tables 5-7, where Table 5 gives the percentage of measured states with an associated optimization cost below the offset, Table 6 gives the optimal cost found, and Table 7 gives the Sharpe ratio of the investment associated with the optimal cost given in Table 6.
We specially focus on the performance of each implementation for the XXL sized DPO problem. In Fig. 4, where we show the optimization cost distribution for each VQE implementation discussed in this section. For reference we include the offset value (dashed gray line) and the Gurobi optimization cost (solid black line). To asses the influence of the quantum noise in this results we include the results of a random sampling in Table 5 and Fig. 4 (gray line). The random sampling reproduces a purely noisy quantum solution. Since our VQE results are obtained without any quantum error mitigation, the random distribution gives us an additional baseline for comparison.
@@
Table 7: Sharpe ratio results of the investment strategies associated to each minimum optimization cost values shown in Table 6. The combinations are cyclic ansatz with CG and RA, ORA and tailored ansatzes with DE, as indicated in the first column, in parenthesis.
@@
Figure 5: Depth and number of parameters for all ansatz considered in section V. Panel a shows the changes in depths (after transpilation for the IBM Torino QPU) for all the ansatzes as we increase the DPO problem size. Panel b shows the increase in the number of parameters of each ansatz for the increasing problem size.
@@ V.1 Cyclic ansatz, Conjugate Gradient optimizer
Following the workflow in Ref. [8], we use the CG for the VQE with this ansatz, which is an iterative method for solving systems of linear equations. Nevertheless, as it strongly relies on gradients being found (which can be a problem due to quantum noise and barren plateaus, as mentioned earlier), it has some limitations when it comes to a large set of variables. Specifically, we find that for all sizes the CG achieves convergence (see Appendix A) to optimization costs very close to the offset of the problem (the average of the random distribution). Thus, we conclude that the standard implementation of the CG is not the best choice for assisting the VQE optimization.
In Fig. 4 (orange line, corresponding to orange line in Fig. 1) we observe that the cyclic ansatz distribution for the XXL DPO problem is close to the random distribution (gray line). This similarity is expected, given the limitations of the CG optimizer and the ansatz depth for this size (depth is 2347, see Fig. 5a, orange dot line). We consider that the optimization reached with this VQE implementation is underwhelming, and in the subsections bellow we explore alternative approaches specifically focusing for this large XXL DPO problem.
@@ V.2 Real Amplitudes, Differential Evolution optimizer
In subsections below, we explore two different methods to further exploit the efficiency of the structure of the RA. Specifically, we use $R_{Y}$-gates and C-NOT-gates connecting first-neighbor qubits. In Subsection V.3, we first modify the RA ansatz to reflect the relations between variables in our problem (represented by qubits), with the main goal of reducing the overall depth of the circuit. Then, in Subsection V.4, we explore an ansatz tailored for the XXL problem, which also focuses on reducing the circuit depth by exploiting both the relations between variables in our problem and the structure of the QPU, aiming to push the resulting distribution further towards optimized values.
@@ V.3 Optimized Real Amplitudes, Differential Evolution optimizer
Next, we propose a systematic circuit structure, the Optimized Real Amplitudes (ORA), based on the RA circuits and the relationship between variables in the QUBO problem. Specifically, we consider RA blocks ranging from the qubits representing the investment $\omega_{a,t}$ of a particular asset $a$ at time $t$ with the qubits representing the investment of the same asset $a$ at time $t+1$, $\omega_{a,t+1}$. This proposal is based on the analysis of the DPO problem formulation in Subsection II.2, where we observe that cross terms only involve investments of different assets at the same time step, or of the same asset for consecutive time steps, $t$ and $t+1$.
@@
This similarity (for the XXL size) can be further appreciated in Fig. 4, where we observe that the ORA optimization cost distribution (solid black line) is very similar to the one found with the regular RA ansatz (solid blue line). The high number of parameters to optimize, combined with the lack of convergence in the ORA optimization (see Appendix A), leads us to adopt an approach aimed to reducing the number of parameters in the next subsection.
@@ V.4 XXL size: Tailored ansatz, Differential Evolution optimizer
We take a step further to design an efficient ansatz circuit tailored for both our problem and the IBM Torino QPU. We consider an intermediate approach that could combine the benefits of a HEA with an ansatz tailored to the QUBO problem [60, 61, 62]. For this purpose, we simultaneously consider the relationship between variables in our QUBO problem, their respective mapping to qubits in the quantum circuits, and the connectivity between qubits in the QPU. This strategy helps minimizing the number of SWAP gates (introduced because of the transpilation) and maximize the parallel execution of gates, thus reducing circuit depth.
@@
The sequence of operations is executed in parallel following the five steps the scheme in Fig. 9b: i) we perform the entire intra-asset circuits, ii) we apply the inter-asset gates, iii) we perform another intra-asset circuits layer, iv) we apply the inter-time gates, and, v) we perform the last intra-asset circuit layer. Note that the choice of this ansatz is not unique. We selected this specific ansatz for its simplicity, but various quantum gates could be used for each transformation (asset, risk, and transaction).
@@
As observed in Table 6, the optimization of this ansatz resulted in a similar lowest cost to the RA and ORA. Additionally, we give in Table 7 the Sharpe ratio (see Eq. 1) obtained for the investment with lowest optimization cost, where this tailored ansatz approach achieves the best Sharpe ratio value.
However, the potential of this ansatz is better appreciated in Fig. 4 (red line, corresponding to red line in Fig. 1), where the optimization cost distribution presents a significant bias towards the lowest optimization costs, and a smaller standard deviation, meaning it has a trend to favor states with optimal associated costs. This confirms the benefits of considering both the QPU and the problem structure when designing the ansatz, enhancing the overall efficiency and accuracy of the quantum algorithm, ensuring that the solution space is explored in a manner aligned with both the hardware capabilities and the inherent features of the problem.

### v2

@@ Abstract
This work explores the potential of the Variational Quantum Eigensolver in solving Dynamic Portfolio Optimization problems surpassing the 100 qubit utility frontier. We systematically analyze how to scale this strategy in complexity and size, from 6 to 112 qubits, by testing different combinations of ansatz and optimizer on a real Quantum Processing Unit. We achieve best results by using a combination of a Differential Evolution classical optimizer and an ansatz circuit tailored to both the problem and the properties of the Quantum Processing Unit.
@@ I Introduction
Figure 1: Optimization cost distributions obtained for the starting VQE approach (orange line), and the fine-tuned VQE implementation (red line) that we develop in this work. These distributions are obtained from QPU calculations in the IBM Torino solving the 112 qubits (XXL sized) DPO problem presented in Section II. For reference we also give the solution obtained using the Gurobi classical optimizer (black dashed line) and the offset of the problem (average of the random distribution, gray dashed line). We add the optimization costs labels “better” and “worse” to reinforce that we address the DPO as a minimization problem. To obtain each cost distribution we account for how many times we found investment trajectories (quantum states) with the same associated cost (rounded to the second decimal value).
@@ II.1 Formulation of the Dynamic Portfolio Optimization problem
$F(\bm{\Omega}),R(\bm{\Omega})$, and $C(\bm{\Omega})$, in Eq. (2) are calculated over a range of $N_{a}$ assets and across $N_{t}$ time steps. In the expression above, $\gamma$ is the risk aversion coefficient. We found that (for the DPO problems solved here) $\gamma=1000$ ensures same order of magnitude for the return and the risk in Eq. (2) [^1] In our work we chose parameter values that ensure all key aspects of the objective function in Eq. (2) are well represented, in order to capture the full complexity of the problem. For example, we do not chose the value of the risk aversion coefficient $\gamma$ according to industry standards, but we chose a $\gamma$ value that ensures the risk function $R(\bm{\Omega})$ and the expected return function $F(\bm{\Omega})$ have the same influence on the final solution.. Specifically, the expected return is defined as
@@
where $\bm{\mu}_{t}=(\mu_{t,0},\dots,\mu_{t,N_{a}-1})$ is the return vector. The components of $\bm{\mu}_{t}$ are given by the logarithmic return  [^2] The expression of the logarithmic expected return we use in Eq. (4) can be traced back to the expected bare return of the investment. Specifically, the $\omega_{t,a}$ money invested in an asset $a$ at time $t$ becomes $\mu^{\text{bare}}_{t,a}\times\omega_{t}$ at the next investment window, with the bare return $\mu^{\text{bare}}_{t,a}=1+(P_{t+1,a}-P_{t,a})/P_{t,a}$. In Eq. (4) we take the logarithmic of the bare return, $\mu_{t,a}=\log(\mu^{\text{bare}}_{t,a})$. of each asset $a$, which are calculated from the changes of their prices at time $t+1$, $P_{t+1,a}$, compared to the previous time price, $P_{t,a}$,
@@
Note that $t=0$ introduces the $\bm{\omega}_{-1}$ term in Eq. (6), which corresponds to the initial condition of the problem. We chose to set $\bm{\omega}_{-1}=\bm{0}_{N_{a}}$ as if there is no initial investment.
@@ II.2 Dynamic Portfolio Optimization problem as a Quadratic Unconstrained Binary Optimization problem
In order to work with the VQE algorithm, we transform the objective function $O(\bm{\Omega})$ in Eq. (2) into a QUBO problem. By definition, a QUBO problem consists of binary variables with terms that are either linear or quadratic, and it must exclude constraints.
@@
$\tilde{\omega}_{t,a}:=\frac{\omega_{t,a}}{K}.$ (10)
@@
Regarding the QUBO formulation, we first tackle the transaction cost function. Due to the absolute value introduced within the transaction costs in Eq. (6), the function is neither quadratic nor linear, so we perform a quadratic approximation of the absolute value, which leads us to
@@
with $x_{t,a,r}$ binary variables, and the subindices $t$ and $a$ representing the time and the asset of the investment, respectively. The subindex $r$ ranges from 00 to $N_{r}-1$, where $N_{r}$ is the amount of binary variables used to describe the normalized investment ($\tilde{\omega}_{t,a}$) for each asset at each time (increasing $N_{r}$ enhances the precision level defining the investment). Note that, with this conversion, the maximum money invested per asset ($K^{\prime}$ in Eq. (11)) is $K^{\prime}=2^{N_{r}}-1$.
Last, we convert the QUBO problem in Eq. (15) into an Ising Hamiltonian, $H_{\text{Ising}}$, since most quantum optimizers (e.g., quantum annealing) focus on finding the minimum eigenvalue of such Hamiltonians (instead of directly operating on the QUBO problem). Thus, we perform a change of variables from the binary variables $x_{t,a,r}$ to Pauli operators $Z_{q}$,
@@ III.1 Classical Benchmarks
Table 3: Sharpe ratio results of the investment trajectories associated to each minimum optimization cost values shown in Table 2.
@@ III.2 Quantum-Inspired Benchmark
In addition to our classical benchmarks, we develop a quantum-inspired Tensor Network algorithm using the Python library QUIMB [39]. We implement a simulated adiabatic evolution (SAE) [40] algorithm, following Ref. [41]. The SAE allows us to have a benchmark for a quantum strategy run on classical hardware. The details on the implementation of the SAE are beyond the scope of the main text and are described on Appendix F.
@@ IV Variational Quantum Eigensolver
We next implement the VQE algorithm to solve the QUBO problem formulated in Eq. (15) in real QPU devices. The VQE algorithm is a leading approach in solving quantum optimization problems, demonstrating significant success on current NISQ-era devices [42, 43, 44]. In this section, we review our implementation of VQE following the scheme in Fig. 3. Specifically, our implementation consist on five steps:
Figure 3: Scheme of the VQE optimization process (described in Section IV): Steps 1-3 comprise generating an ansatz circuit for a given set of parameters $\bm{\theta}^{(i)}$ (step 1), measuring the expected value of the Ising Hamiltonian $H_{\text{Ising}}$ using the estimator primitive of IBM quantum services (step 2), and optimizing $\bm{\theta}^{(i)}$ with a classical computer (step 3). If the optimal $\bm{\theta}^{(i)}$ is not found in step 3, the process returns to step 1. Once the optimal $\bm{\theta}^{(f)}$ is found, the algorithm proceeds to sample the solution (steps 4 and 5), which involves sampling the ansatz circuit for the optimal $\bm{\theta}^{(f)}$ (step 4, where we use the sampler primitive of IBM quantum services) and identifying the optimal sampled state (step 5).
@@
    Finally, we compute the optimization cost associated with each bit string, $c(b)=\braket{b}{\hat{H}}{b}$, and identify the solution bit string |$\ket{b}$ resulting in the smallest $c(b)$ value.
@@ V Quantum Dynamic Portfolio Optimization
As mentioned in the previous section, the performance of the VQE algorithm strongly relies on the chosen classical optimizer and ansatz. In this section, we explore this dependence as we scale the DPO problem from a size requiring 6 qubits to a large scale problem requiring 112 qubits (see Table 1) and we systematically solve each problem size for a better comprehension of the scaling for each method. Upscaling the problem not only increases the number of qubits but also the depth of the circuits, leading to decoherence and gate-associated quantum noise. We remark that in our work we do not use any quantum error mitigation. Thus, we aim to find VQE implementations that are robust enough to quantum noise such that their raw implementation gives satisfactory results.
@@
It is important to note that every circuit must be transpiled [49] before running it on QPU. This transpilation primarily depends on the native gates of the quantum computer and the arrangement of its physical qubits. For example, if two logical qubits that need to be connected are not directly linked on the QPU, SWAP gates are introduced to achieve this entanglement. The transpilation process can significantly increase the depth of each circuit. In this work, we discuss depth after transpilation.
The comparison between methods for all problem sizes in the subsections below is built upon the results in Tables 5-7, where Table 5 gives the percentage of measured states with an associated optimization cost below the offset, Table 6 gives the optimal cost found, and Table 7 gives the Sharpe ratio of the investment associated with the optimal cost given in Table 6.
We specially focus on the performance of each implementation for the XXL sized DPO problem. In Fig. 4, we show the optimization cost distribution for each VQE implementation discussed in this section. For reference we include the offset value (dashed gray line) and the Gurobi optimization cost (dashed black line). To asses the influence of the quantum noise in this results we include the results of a random sampling in Table 5 and Fig. 4 (gray line). The random sampling reproduces a purely noisy quantum solution. Since our VQE results are obtained without any quantum error mitigation, the random distribution gives us an additional baseline for comparison.
@@
Table 7: Sharpe ratio results of the investment trajectories associated to each minimum optimization cost values shown in Table 6. The combinations are cyclic ansatz with CG and RA, ORA and tailored ansatzes with DE, as indicated in the first column, in parenthesis.
@@
Figure 5: Depth and number of parameters for all ansatz considered in section V. Panel a) shows the changes in depths (after transpilation for the IBM Torino QPU) for all the ansatzes as we increase the DPO problem size. Panel b) shows the increase in the number of parameters of each ansatz for the increasing problem size.
@@ V.1 Cyclic ansatz, Conjugate Gradient optimizer
Following the workflow in Ref. [8], we use the CG for the VQE with this ansatz, which is an iterative method for solving systems of linear equations. Nevertheless, as it strongly relies on gradients being found (which can be a problem due to quantum noise and barren plateaus, as mentioned earlier), it has some limitations when it comes to a large set of variables. Specifically, we find that for all sizes the CG achieves convergence (see Appendix A) towards optimization costs very close to the offset of the problem (the average of the random distribution). Thus, we conclude that the standard implementation of the CG is not the best choice for assisting the VQE optimization.
In Fig. 4 (orange line, corresponding to orange line in Fig. 1) we observe that the cyclic ansatz distribution for the XXL DPO problem is close to the random distribution (gray line). This similarity is expected, given the limitations of the CG optimizer and the ansatz depth for this size (depth is 2347, see Fig. 5a, orange dot line). We consider that the optimization reached with this VQE implementation is underwhelming, and in the subsections below we explore alternative approaches specifically focusing for this large XXL DPO problem.
@@ V.2 Real Amplitudes, Differential Evolution optimizer
In subsections below, we explore two different methods to further exploit the efficiency of the structure of the RA. Specifically, we use $R_{Y}$ gates and C-NOT gates connecting first-neighbor qubits. In Subsection V.3, we first modify the RA ansatz to reflect the relations between variables in our problem (represented by qubits), with the main goal of reducing the overall depth of the circuit. Then, in Subsection V.4, we explore an ansatz tailored for the XXL problem, which also focuses on reducing the circuit depth by exploiting both the relations between variables in our problem and the structure of the QPU, aiming to push the resulting distribution further towards optimized values.
@@ V.3 Optimized Real Amplitudes, Differential Evolution optimizer
Next, we propose a systematic circuit structure, the Optimized Real Amplitudes (ORA), based on the RA circuits and the relationship between variables in the QUBO problem. Specifically, we consider RA blocks ranging from the qubits representing the investment $\omega_{t,a}$ of a particular asset $a$ at time $t$ with the qubits representing the investment of the same asset $a$ at time $t+1$, $\omega_{t+1,a}$. This proposal is based on the analysis of the DPO problem formulation in Subsection II.2, where we observe that cross terms only involve investments of different assets at the same time step, or of the same asset for consecutive time steps, $t$ and $t+1$.
@@
This similarity (for the XXL size) can be further appreciated in Fig. 4, where we observe that the ORA optimization cost distribution (solid black line) is very similar to the one found with the regular RA ansatz (solid blue line). The high number of parameters to optimize (and lack of convergence) of the ORA approach (see Appendix A), leads us to adopt an approach aimed to reducing the number of parameters in the next subsection.
@@ V.4 XXL size: Tailored ansatz, Differential Evolution optimizer
We take a step further to design an efficient ansatz circuit tailored for both our problem and the IBM Torino QPU. We consider an intermediate approach that could combine the benefits of a HEA with an ansatz tailored to the QUBO problem [60, 61, 62]. For this purpose, we simultaneously consider the relationship between variables in our QUBO problem, their respective mapping to qubits in the quantum circuits, and the connectivity between qubits in the QPU. This strategy helps minimizing the number of SWAP gates (introduced because of the transpilation) and maximizing the parallel execution of gates, thus reducing circuit depth.
@@
The sequence of operations is executed in parallel following the five steps the scheme in Fig. 9b: i) we perform all intra-asset circuits, ii) we apply the inter-asset gates, iii) we perform another intra-asset circuits layer, iv) we apply the inter-time C-NOT gates, and, v) we perform the last intra-asset circuit layer. Note that the choice of this ansatz is not unique. We selected this specific ansatz for its simplicity, but various quantum gates could be used for each transformation (asset, risk, and transaction).
@@
As observed in Table 6, the optimization of this ansatz resulted in a similar cost to the RA and ORA. Additionally, we give in Table 7 the Sharpe ratio (see Eq. 1) obtained for the investment with lowest optimization cost, where this tailored ansatz approach achieves the best Sharpe ratio value.
However, the potential of this ansatz is better appreciated in Fig. 4 (red line, corresponding to red line in Fig. 1), where the optimization cost distribution presents a significant bias towards the lowest optimization costs, and a smaller standard deviation, meaning it has a trend to favor states (investment trajectories) with optimal associated costs. This confirms the benefits of considering both the QPU and the problem structure when designing the ansatz, enhancing the overall efficiency and accuracy of the quantum algorithm, ensuring that the solution space is explored in a manner aligned with both the hardware capabilities and the inherent features of the problem.